A force is applied to a 2.6 kg mass and produces 5.9 m/s[sup]2[/sup] acceleration.What acceleration would be produced by
the same force applied to a 14.9 kg mass?
1
Expert's answer
2010-11-08T02:32:55-0500
Firstly let's find the force using the second Newton's law:
F = ma = 2.6kg * 5.9 m/s2 = 15.34 N.
Acceleration would be:
a = F/m' = 15.34 / 14.9 = 1.03 m/s2